On a trip, are 5/9 of the way to destination after 2 1/2 hours. You are traveling at a constant speed. When will you reach your destination?

Answer by ankor@dixie-net.com(22740)

On a trip, are 5/9 of the way to destination after 2 1/2 hours. You are traveling at a constant speed. When will you reach your destination?
:
Let t = time required to complete the trip:
:
(5/9)t = 2.5
:
Mult both sides by 9 and you have:
5t = 9(2.5)
:
5t = 22.5
:
t = 22.5/5
:
t = 4.5 hrs to complete the trip
:
:
Check on a calc, enter (5/9) * 4.5 = 2.5
:
Not that hard, right?